This page: http://monitor.mikes.fi/ionclock/

2019 archive

  • 2019-11-08: -C2 Zeeman component linewidth estimate of 800 Hz.
  • 2019-11-05: Zeeman peaks with B-shield (1.6 uT estimate).
  • 2019-11-04: Now with first B-shield. Laminated plywood + metglas.
  • 2019-07-12: Zeeman peaks in ambient (66uT estimate) B-field.
  • 2019-07-10: Initial clock-transition search...
  • 2019-07-10: Page created.

  • 2019-12-20

    08:00Z ion seems to have disappeared!?
    Try a bit of 405, 1033, and 461..., and re-lock to Rb-line.

    2019-12-19

    10:00Z Longer run over lunch, 6x6x6 = 216 pts. n_phot=300k
    11:30Z progress to pt 93/215 - so this run is >2 hours long..
    Optimal voltages [ 0.24380421 26.67929547 43.62223335]
    Data file 5009
    Contrasts at Vopt ( 2.5, 2.7, 2.8) %

    Ball-size is abs(contrast), +sign is green, -sign is red. sum figure shows sqrt(A^2+B^2+C^2)
    optimal point is blue ball (size not to scale) in sum-figure.


    0900Z: Run the NPLv3-umotion-optimizer, around previous voltages, with n_phot=100k
    4x4x4 = 64 voltage-combinations, x3 beams. Data file 5004.
    New optimal voltages (-5.5 V12cm): [ 0.24315127 26.67666472 43.61507129]
    V1 = -5.7432, V2 = -5.2568, V3 = -16.938399999999998, V4 = 70.2918
    Here n_phot=300k contrasts are: (2.5, 2.7, 3.1) % - (?not so great - same as before?)

    micromotion contrasts are now slightly worse.
    (A, B, C ) = (2.4, 3.1, 4.8) % repeated (2.0, 3.1, 4.6) % - measured with n_photons=300k and n_bins=25
    voltages from y-day were (-5.5, +0.2668, +26.719, +43.503) V = (V12cm, V12dif, V34cm, V34dif)

    Fluorescence from all beams ABC was quite steady at 37(1) kCPS through the night.
    This is about 700 CPS in the usual 20ms detection window

    2019-12-18

    at some point Rb-lock was lost! detuning was down to 170MHz for tickler - this made the lock jump to the next Shiner-peak?
    when locked to wrong shiner-peak, fluorescence was 200counts/20ms (required long camera exposure) - after re-locking to right peak >700counts/20ms.

    micromotion investigations. How does V12cm change the optimal u-motion voltages?
    d V12dif / d V12cm = -0.81 mV/V
    d V34cm / d V12cm = 1.30 V/V - strong coupling!?
    d V34dif / dV12cm = 18 mV/V

    V12cm controls the secular ratio
    V12cm = 0 V -> wz/wr = 1.52 and 1.57
    V12cm = -3.5 V -> wz/wr = 1.76 and 1.84
    V12cm = -5.5 V -> wz/wr = 1.93 and 2.02

    secular frequencies with tickler
    170MHz detuning seems to work (but see note on Rb-lock!), see rise in fluorescence at resonance
    DDS attenuation 4dB for radial modes, 10 dB for z-mode
    At -5.5 V12cm, RF 70mVpp, seculars (0.936125, 0.9805, 1.8918) MHz. Radial modes split by 44 kHz

    video: https://www.youtube.com/watch?v=3vQdb2Ptwc8

    overnight fluorescence monitor, gate time 1000ms, n_pts = 46800 (13 hours)
    NOTE all beams ABC on!
    data: 2019-12-18-counts.txt

    2019-12-17

    256x speedup stream: https://www.youtube.com/watch?v=bSQgo7BL8f4
    slow stream: https://www.youtube.com/watch?v=Ey-IkPURY4o

    overnight scans, variations in probe length and attenuation.
    5kHz span, 10 Hz resolution, 499 pts/scan, ca 1 hour/scan
    'center' scans, not randomized (seems to scan outwards from center?)
    scan1: 20 dB, 5 ms
    scan2: 20 dB, 10 ms
    scan3: 20 dB, 20 ms
    scan4: 20 dB, 40 ms
    scan5: 22 dB, 10 ms
    scan6: 22 dB, 20 ms
    scan7: 22 dB, 40 ms
    scan8: 22 dB, 80 ms
    scan9: 18 dB, 10 ms
    scan10: 18 dB, 20 ms
    scan11: 18 dB, 40 ms
    scan12: 18 dB, 80 ms
    scan13: 18 dB, 5 ms


    run5: same scan as run4, longer 20ms probe
    run6: longer 40ms probe


    run4: try to zoom in on left peak (-C1?)
    center 75.9474 MHz, span 1 kHz, resolution 10 Hz, 101 points. Probe 10ms, attenuation 20 dB


    run3: try I3=0mA, to separate Zeeman-components? Coil currents (-5, 0, 0) mA
    center 75.9484 MHz, span 4 kHz, resolution 100 Hz, 41 points. DDS att 20 dB (from rabi-osc attempts 20 dB seems fine?)


    run2: center 75.9484 MHz, narrower span 4 kHz, lower resolution 25 Hz, 161 points, DDS att 5 dB, 10ms probe, 20ms detection, 500/50 bright/dark.


    run1: center 75.948 MHz, span 12 kHz, resolution 100Hz, 121 points. DDS att 5 dB. 10ms probe. 20ms detection 500/50 bright/dark.
    ca 7s per 100 pulses
    record-high 48% excitation!?
    B-field seems low, peaks overlap


    f_beat = 23.150
    estimated f_corrected -330.396 MHz

    Micromotion
    NPL-optimization gives (0.2668, 26.719, 43.503) voltages. Here contrasts are (0.3, 0.5, 0.7)%. n_photons=300k

    With no changes to bias-voltages, get contrasts of (5.5, 8.7, 13.3) %, detuning 210 MHz
    The old voltages were (?) (0.246, 26.57, 44.2) V

    2019-12-17 13:40Z Loading 1.60A, 6min 20s

    setup changes: metglas 3rd layer. Toptica clock-laser (lock servo might not be optimal)

    2019-12-10

    Simulation: how many photons to collect into the correlation histogram?
    <1% stdev in the S-contrast requires more than 2e4 photons.
    ca 0.5% stdev in S expected with 1e5 photons.
    <0.1% stdev requires >1e6 photons
    Number of bins in the histogram is not important

    2019-12-07

    15:00Z: ion was dark from 23-local y-day evening. 405 or 422-re-lock doesn't help. shut down lasers.

    2019-12-06 Independence Ion

    Stream: https://www.youtube.com/watch?v=z9VFbs4FogY
    128x speedup: https://www.youtube.com/watch?v=f3edhdwqXgc

    scan1 Lorentzian fits to data, based on f_corr estimate -330.3895 MHz
    lorentzian(x) = height*(width)/(pow(x - center,2) +pow(width,2))
    Estimated AHM3 vs UTC is -2.5194e-12 or 1121 Hz @ SRS
    Componentheightwidth / Hzcenter-SRS / Hz
    -C42.1e-3241-9072
    +C42.7e-3225+11291
    C4 pair--+1110
    -C21.5e-3182-2949
    +C21.7e-3200+5179
    C2 pair--+1115
    -C12.3e-3203-937
    +C11.9e-3203+3114
    C1 pair--+1089

    Estimate of B-field based on scan1 pairs is ~0.363(1) uT
    Pairpeak-separation / HzZeeman shift / Hz/uTB = sep/(2*shift) / uT
    C140515598.50.362
    C28128111970.363
    C42036327992.50.364

    from livestream, ion lost 20:59:30Z. It was loaded 2019-12-04 13:37Z.

    leave longer ca 1h scans running
    center 75.9455, range 12 kHz, resolution 25 Hz, 481 points
    probe 10ms, attenuation 20 dB
    seem to detect C1 through C4 (not C5)
    B ~0.36 uT with I = (-5, 0, -1) mA applied

    data file e.g. "2019-12-06-scan1.txt"
    scan1 (4794) through scan7 (4800) contain good data. scan8 (4801) has both a 5min dark period and the ion loss event.
    scan9 through scan24 are invalid data

    run5


    12:15Z suddenly lost ion??... and then it returned at 12:20z after 5 min!?

    run4
    try I3 = 3 mA
    -C1 at 75.9412
    +C1 at ?
    Run with a few different currents and plot in same figure
    orange (0, 0, +3) run4: I3 = +3 mA
    red (0, 0, +1) run4-1mA: I3 = +1 mA
    pink (0, 0, -1) run4-minus1mA: I3 = -1 mA
    blue (I1, I2, I3) = (+1, 0, -1) (did not move -C1 closer to line-center)
    green (-1, 0 , -1) - now closer!
    cyan (-2, 0, -1) - best so far
    yellow (-3, 0, -1) - we now se also -C2 in the 3kHz scan? excitation-P for -C1 is lower? ex-P for -C2 is record-high!? run also the +side
    tab:blue (-4, 0, -1)
    tab:brown (-5, 0, 1)


    run3
    change I-driver state from "#" (edit mode?) to "^" (actual current-drive?) - maybe this will move C1?
    same centers and ranges as before
    +C1 at 75.94895 (moved only -100Hz?)
    -C1 at 75.9415 (did not move much)


    run2
    search wider for C1
    old (1mA) predicted center at 75.9416 (-C1) and 75.9486 (+C1)
    observed at 75.94145 (25 jumps) and 75.94905 (26 jumps)
    seems like the 2mA value is not in use? (repeat?)


    run1, I3=2mA
    center 75.9451 (no change)
    range 7 kHz (should see C1 peaks, they were at +/-3.5 kHz previously)
    step 100 Hz, 71 points


    Increase I3=2mA - prediction is that this further decreases B.
    interrupt run8 from y-day, it is taking too long...

    2019-12-05

    Clock transition run

    run8
    longer run, leave overnight
    win-machine spontaneously rebooted in the evening (~21 local) !?! live-stream cut


    4753 26dB attenuation rabi-flop
    4752 rabi-flopping on -C2, DDS att 23 dB, range 1ms to 20ms

    run7
    zoom in on -C2 and get a linewidth (with first B-shield 2019-11-05 linewidth was ~800Hz )
    Zeeman model predicts -C2 at AOM=75.9381 MHz


    run6
    peaks fit a B-field of 1.25 uT (but the prediction was 0.2uT from 1mA of current?)
    try to get a wider view
    back to calculated center (+2800Hz from run5)
    wider range 30kHz, resolution 200Hz, probe 10ms, 151 points
    Zeeman peaks offset +1.1 kHz from scan-center, with B=1.25 uT


    run5
    add 1mA of I3 coil current. ca 200uT/A -> B=0.2 uT
    This should split C5 by 7.8kHz and C1 by 1.12kHz
    larger range 10kHz but lower resolution 100Hz -> 101 points
    20ms probe


    run4
    double range to 1kHz. 40ms probe, more attenuation 20 dB (seems OK from rabi-flopping)
    10Hz resolution, 101 points


    Now see high excitation probabilities, around 30% at best.
    (4743) Rabi-flop, DDS_att = 5dB. Rises from 1ms to 4ms - then flat
    (4744) Rabi-flop 10 dB att, rises until 8ms then flat
    (4745) Rabi-flop 20 dB att, intially quadratic - good(!?) - from 0 to 6ms
    (4746) Rabi-flop 30 dB att, narrower range 1ms to 30ms. too much attenuation - not much excitation.
    (4747) 23 dB att.

    run3
    shift down center -2800 Hz
    longer 40ms probe, 10 Hz resolution 1 kHz range
    51 points


    run2
    longer probe 20ms, lower resolution 25Hz, larger range 2*6 kHz. 241 points. N_probe=100

    run1
    tau_probe=10ms, DDS_att=5dB, N_pulses=100, tau_det=20ms, N_bright=500, N_dark=50

    6th run of NPL, over lunch, gave (0.258, 26.594, 43.829)
    contrasts(300 kphotons, 35 bins) are (A, B, C) = (0.94%, 1.3%, 0.79%)
    300 kphoton measurement takes (16, 23, 25) s to perform, depending on fluorescence amount from each beam

    NPL-algorithm with 3 variables now seems to converge
    detuning AOM 210MHz, V12cm = -5.5V (NOTE: this is FIXED in the v2 NPL-script!)
    ABC-optimal point, 1st run: (0.250 , 26.573 , 43.814 ) (V12dif, V34cm, V34dif)
    Here ABC-contrasts (2, 2, 2) %
    60 k photons/histogram, 3 points scanned per voltage (27 points)

    ABC-optimal point, 2nd run: small changes to optimum (-1 mV, +5 mV, +8 mV)
    voltages = (0.249, 26.578, 43.822)
    100 k photons/histogram, 5 points per voltage (125 points total)
    here ABC-contrasts (2.5, 2.5, 2.5) % - not any better than previous optimum

    2019-12-04


    NPL-algorithm modified to only have 3 scanned parameters (V12dif, V34cm, V34dif)


    Dashed lines show A-optimal and B-optimal voltages. Dash-dotted lines are shifted versions (2020-04) - probably not useful
    pa = -1.5009 85.1762
    pb = 1.6786 2.3214
    For example a change of (V34cm, V34diff) = (+1, -1.5) should be along the A-optimal line.
    These should be compared to the planes found in the 2019-12-19 data

    AB-contrast optimal point found in (V34cm, V34dif) plane. C-contrast not optimal.
    V-shaped bias-sweep plots, sine-shaped correlation histograms.

    Micromotion minimization, need to move both ABC beams and ASE as bias-voltages change
    Using V12cm = -5.5V to get axial/radial ~2

    13:37Z Ion loaded 1.60A / 1.25 W and 6min 35s

    2019-12-03

    Filter-boards removed from clock-laser prestabilizer - now seems to stay locked for >8h.
    Cavity drift from 8h of data: +18.093 mHz/s
    OHDEV(drift-corrected) reaches 1e-15 at around 200s.
    f_cavity = -490.2075 MHz (SRS) (morning 2019-12-03)

    2012-12-02

    Ion loaded 1456Z (1.60A and 8min 10s) - but lost 22:xxZ the same day

    By scanning, found voltages that minimize A-contrast (but B/C beams do not cool in this state?)
    Vbias = (-5.5, +1, 0, 80.4) V
    A-contrast (repeats, n_photon = 60k or 80k): 0.9%, 0.45%, 1.2%, 0.58%
    Clear minimum with pi-phase change at V34diff 80.4V. Sine-shaped correlation histogram

    Secular frequencies: could not find radial splitting, see only one peak. why? (using A-beam)
    70mVpp from agilent, and the usual 1.65Vpp monitor.
    ARTIQ DDS as tickler, 4 dB attenuation for Radial, 10 dB attenuation for Axial.
    Fluorescence dips or peaks, depending on cooling laser detuning.
    Siglent siggen+bias-amp connected for DC-modulation (try this later also with shorting N-connector on helical-resonator!)
    V12common / V Axial/Radial ratio Axial / MHz Radial / MHz
    0 1.58 1.72 1.09
    -2 1.72 1.7905 1.0387
    -4 1.88 1.854 0.9846
    -5 1.97 1.8848 0.9563
    -6 2.06 1.9142 0.927

    abs(qz)=0.32, trap voltage V0 = 390 V (assumes 0.75 trap efficiency)
    az = 0 at V12common = -5.319 V

    Loading at 1.70A gave two (or more?) ions

    2019-11-29

    Cavity modes
    Estimated cavity mode spacing 501.305 MHz
    Note that comb measures drift-corrected IR, not cavity frequency.
    f_cavity - SRS / MHz sign, f_beat / MHz Wavemeter / nm Comb-line nr Cabity mode nr
    -490.213 -23.145 1348.052 2 223 894 887 242
    +11.092 +27.508 1348.051 2 223 896 887 243
    +512.396 -21.840 1348.049 2 223 899 887 244
    If cavity mode-numbers and abs-frequencies correct, mode spacing is 501 304 565 Hz
    (measured against AHM3, with estimated frequency of -2.51e-12 wrt UTC(MIKE), maser is slow)

    2019-11-28

    Observed drift in 4h run: +24.733 mHz/s

    Increase drift-correction by +4.49 mHz/s (single-pass), based on 4-hour run 2019-11-28 evening.

    2019-11-25/26


    ion lost 2019-11-26 at 17:24:52Z, storage time was ca 29 hours

    Clock laser (briefly) locked to cavity
    F_corrected = -330.386205 MHz
    f_beat = 23.145390(50) MHz
    With fixed 178.5 MHz AOM this gives line-center at AOM_Zeeman = 75.943103 MHz P_674_A-beam = 44 nW (5dB att. on Zeeman AOM)
    20 ms detection window, bright=230, dark=30

    Drift-correction was not updated recently, still runs with AOM-drifts of k0 + k1 = (6.532e-3 + 2.728e-3) Hz/s = 0.5 * 18.52 mHz/s

    Try to optimize micromotion - but no clear minimum found
    Longer NPL-algorithm run gives V12diff, V34common, V34diff = (+2.22, +23.138, +105.991)
    This gives photon-correlation contrasts A/B/C = 18% 4% 14%
    About 60 000 photons in one histogram seems necessary for good statistics

    New RF-drive 16.716 MHz. 70mVpp gives 1.65 Vpp monitor as before.
    2019-11-25 12:21Z Ion loaded with I_disp = 1.60 A after 4min 45s.

    Bias-voltage source improved range to +/-168 V on V3 and V4

    2019-11-14

    Helical-resonator shield-adjustments.
    After tuning impedance match (New RigExpert AA-55 Zoom)
    - SWR 1.03 @ 16.7188 MHz
    - Z=48.9 Ohm, R= 48.9, X= -0.96
    - Q ~730...770

    2019-11-08

    pulse data from wide scan of 2019-11-05
    30ms
    26ms

    Overnight results in Technicolor 3D:


    (07:38) ion lost, tried 405nm but no change, checked 422nm Rb-lines - seems OK
    from video, lost around 2019-11-07 21:39Z (was trapped at 2019-11-04 09:52Z)
    camera intensifier battery-box needs a new 9V battery..

    (07:33Z) clock-laser has unlocked at some point. when?

    2019-11-07

    Observed line is ca.800Hz (optical) wide - what B-field changes does this correspond to?
    C2 sensitivity 11.190 kHz/uT, so 400Hz shift by 0.036 uT.
    this is 2% of the assumed 1.56uT field at the ion
    If we have a 36 nT peak AC-field (frequency?), can this be measured with a probe+meter?
    How? Outside the shield, close to ion, and assume some AC attenuation?

    (14:55Z) new center for scan: 75.93805 MHz
    schedule overnight runs, 8pcs a 2hours = 16 hours.
    all scans have same AOM range 1200Hz and AOM step 5Hz, 239 frequency-points per scan.
    3rd run 50ms
    50ms_run3
    5th run 40ms
    40ms_run5
    1st run 30ms
    30ms_run1
    7th run 20ms
    20ms_run7
    2nd run 10ms
    10ms_run2
    4th run 5ms
    5ms_run4
    8th run 2.5ms
    2.5ms_run8
    6th run 1ms
    1ms_run6

    (14:40Z) New drift-compensation = [6.523 (old) + 2.728 (new) ]*2 = 18.52 mHz/s, from 58794.611

    (13:17Z) combined plot, run1 and run2


    (13:01Z) What effect does cavity-drift have? We currently compensate 2*6.532 mHz/s = 13 mHz/s
    Measured overnight drift is 18.5 mHz/s
    During 1h we have uncompensated drift of 3600s*5 mHz/s = 20 Hz
    During 1 day, 470 Hz!

    (11:54Z) start run2: same 60ms probe, range 1500Hz (down from 2kHz), AOM step 5 Hz (down from 10Hz). 200 shots/freq.
    301 points, 25s/pt, expected duration 2h, will finish at ca.13:54Z
    run2

    re-lock clock laser
    (values are 15 min averages, from Chronograf)
    f_beat= 23.148089 (37) MHz (12:01Z)
    f_cavity = -490.255964 MHz
    f_corrected = -330.391669 MHz
    cavity_drift = 0.024 Hz/s

    Clock-laser unlocked during lunch-break scan...(11:21UTC)

    Use only A beam, P_674 = 105 nW, try 200 pulses per frequency
    Same -C2 center-frequency (75.9383 MHz), 2kHz range (down from 7kHz), 10 Hz resolution (down from 20Hz).
    200 frequency points, 25-26s/point, duration 1.5 hours
    run1 60ms: run1

    overnight cavity frequency: 2019-11-07_fcavity.csv

    2019-11-06

    NOTE: all three beams ABC ON for cooling and probe
    Scan area around -C2, with different probe pulse lengths= 30ms, 40ms, 50ms, 60ms, 70ms....
    7200 Hz AOM scan range, 20Hz AOM resolution, 360 points per scan. (1 hour per scan?)
    shorter detection window 10ms, collects ~350 photons in 10 ms.
    -C2_fcenter = 75.9383 MHz
    Fcavity = -490.257406 MHz
    Fcorrected = -330.392137 MHz
    f_beat = 23.148357 MHz
    clock-cycle configuration (clearout, cool, detect, repum)=(10, 50, 10, 60)
    bright=350, dark=25. threshold is geometric mean
    run1 30ms: run1 (note randomized order of frequencies!)
    run2 40ms: run2
    run3 50ms: run3
    run4 60ms: run4
    run5 70ms: run5
    run6 30ms: run6
    run7 40ms: run7
    run8 50ms: run8
    run9 60ms: run9
    run10 70ms: run10
    run11 80ms: run11
    run12 90ms: run12
    run13 30ms: run13
    run14 40ms: run14
    run15 20ms: run15

    all:


    Clock laser unlocked at some point during morning? (was unlocked at 0740 UTC)
    Some of the later data in the overnight run is bad.

    2019-11-05

    overnight runs

    Overnight scheduled runs for 30, 26, 22, 18, 14, 9, 6, 3 ms probe pulses. Approx 10 nW in A-beam to ion.
    Scan is 75.947MHz +/- 35 kHz (assumes B=50uT/b_att=1.56 uT is correct...?).
    1000 points, approx 70 Hz AOM resolution, 140 Hz optical.

    30ms lin scale


    22ms lin sale


    all scans, log scale:

    Zeeman notes


    Zeeman-split at 50uT is +/- 281, 560, 1121, 1400, 1961 kHz (for C1, C2, C3, C4, C5 components)
    sensitivities 5.6124, 11.190, 22.415, 27.992, 39.217 kHz/uT (NRC email 26.5.2013)
    Corresponding to gS=2.00248 gD=1.200486 and gS/gD = 1.66806
    Barwood2012b reports gS/gD = 1.668057
    Normalized by C1 sensitivity ratios are approx: (1, 2, 4, 5, 7) and differ from integers by about 0.02 at most.
    Ratios for zeeman-split from above sensitivities are (1, 1.9938, 3.9938, 4.9875, 6.9876)

    clocklaser notes

    run1

    try minimizing micromotion, without much success... need higher Vbias voltages?
    lower 674 intensity 10-fold, to around 10nW at ion (28dB attenuation on 178.5MHz DDS AOM, ca 2uW at 674-launch).
    use 3ms probe pulse
    scan range around inner two Zeeman-components (from -C2 to +C2) from 2019-11-04.
    1000 frequency points (ca 100Hz AOM-step), ca 2 hours.

    Fit to data!?
    center_freq = 75.947000 (+2 kHz SRS)
    C5-shift = 0.030641
    +C5 = 75.977641
    -C5 = 75.916359
    B_eff = 50uT/b_att = 1.56 uT

    2019-11-04

    2019-11-04 run2:
    from 75.85 to 76.05 (200 kHz) in 5000 steps (~10 hours)
    40 Hz AOM step, 80 Hz optical
    try 16 ms probe pulse (?)


    2019-11-04 run1: from 75.8465 to 76.0465 (200 kHz) in 1000 steps (~2 hours)
    200Hz AOM step, 400 Hz optical step
    first half of data is bad, ASE-overrides were ON...
    8 ms probe pulse, maybe 100nW in A-beam.

    2019-10-31 photo

    Cooling laser detuning scan
    Micromotion minimization with the Baynham-matrix was successful (?).
    Measured contrasts before scans (A/B/C) = (1.0 % / 1.6 % / 5.3 %)
    However only A and B beam scans were successful, data from C-beam looks strange?
    Measured power (A/B/C) = (0.38 / 0.35 / 0.27) uW
    Theory curve has detuning shifted by -1 MHz, background counts 1700 cps estimated, fluorescence collection efficiency 1/218

    2019-10-30 photo

    Canon EOS 60D, 5 s exposure, f/2.8 105mm Sigma Macro, ISO-3200. 13mm extension tube.

    2019-07-12 Weekend

    The 8ms results were interpreted as due to either 33uT or 66uT ambient B-field.


    2019-07.13 ca 17:00+03 re-locked cooling laser - strange noise on Rb-signal!?
    8ms pulses: sat-13-8ms.txt

    Ion trapped on wednesday (after lunch?) was lost friday 19:50+03. 405nm did not help (H-Sr molecule??), A/B/C beams did not help.
    Better diagnostics for cooling-laser needed - not sure if servo was at limit or not..
    New ion trapped sat 2019-07-13 ca 10:30+03
    New longer probe-time runs scheduled
    2.0 ms, 4.0 ms, 8.0 ms

    8000 frequency-points in 2 MHz span -> AOM step 250 Hz -> optical step 500 Hz.
    100 pulses per frequency, ca 7 s per frequency or 16 hours per 8000 frequency-points.
    674nm ca 100nW, 422nm ca 1.1 uW detuning_AOM=200MHz, beam A.
    run1: probe 0.5 ms
    run2: probe 1.0 ms
    run3: probe 2.0 ms
    run4: probe 4.0 ms

    2019-07-12 Friday

    1ms, lower power

    Zeeman AOM att. 21 dB (was 15dB) - now approx 50 nW to ion

    4ms

    longer 4 ms pulse
    A beam, ca 100 nW

    2ms

    longer 2ms pulse
    A beam, ca 100 nW

    2019-07-11 night

    one beam, the A beam
    cooling ca 1.5 uW
    clock ca 100nW, 1ms pulses
    detection window longer (was 20 ms), now 34 ms
    AOM-scan-batch is 250 kHz / 1001 points. 8 batches = 8008 points in total

    2019-07-11 day

    2019-07-10 run

    Optical frequencies denoted SRS are offsets from BIPM SRS value.
    All three cooling beams, AOM = 200 MHz detuning.
    All three clock beams, ca 100 nW in total, 1 ms pulse.
    Estimate of drift-corrected cavity-light: -330.486522 MHz SRS.
    AOM-shift from this is AOM_fixed +178.5 MHz + 2*AOM_Zeeman
    So the high peaks are centered at +0.854 MHz SRS and split by +/-0.617 MHz
    Jumps (more than 2/100):

    2019-06-18

    Ion photo
    Canon 60D, Sigma f/2.8 105mm macro + XX mm extension tube, ISO-1000, 8 s exposure



    Secular frequencies measured with tickler (on one outer electrode)
    Here -6 VDC is applied to both outer electrodes in order to get an axial/radial ratio of ~2.005.
    The common mode lowers the radial frequency and raises the axial frequency
    Radial splitting ~43 kHz.

    Without common-mode voltage on the outer electrodes the axial/radial ratio is ~1.54.
    Radial splitting ~37 kHz.